# Learning county from pixels: Corn yield prediction with attention-weighted multiple instance learning

### v3

@@ 2 Data acquisition
Category Variables Unit Related Properties Spatial Resolution Source Latency
@@
Soil Organic Matter (SOM) $\mathrm{kg}$/$\mathrm{m}^{2}$ Soil nutrient uptake
Cation Exchange Capacity (CEC) $\mathrm{cmol}$/$\mathrm{kg}$
@@ 3.1 Problem definition
To formalize the county level crop yield prediction problem, let $n$ be the number of counties in the given study area. Let $I_{i}$ be the satellite imagery and $y_{i}$ be the yield for each county $i$, $i=1,...,n$. Each image is on the space $\mathbb{R}^{W\times H\times C}$, where $W$ and $H$ are the width and height of the image, and $C$ denotes the channels. We calculate VIs from the satellite imagery, along with other agricultural and environmental features (weather, soil properties, etc.), to form our county-level input feature $F_{i}$. From a machine learning perspective, our training objective is to develop a model $f$, $f(F_{i})=y_{i}$.
@@ 3.2 Multiple instance learning
Problem formulation for MIL in regression task For the county-level corn yield prediction task, we have the yield (label) of each county (bag), and each county is composed of a number of pixels (instances). By utilizing the detailed information of each county, MIL can improve prediction accuracy. Suppose there are $m_{i}$ pixels in each county $i$, $i=1,...,n$. To differentiate from the county-level feature $F$, we use $P$ to represent pixel-level features. For each pixel $j$, $j=1,...,m_{i}$, we extract its feature $P^{j}_{i}$. There exists a model $f$ with many-to-one mapping, namely $f(P^{1}_{i},P^{2}_{i},...\ldots,P^{m_{i}}_{i})=y_{i}$. We hope our model can automatically learn the structure of a county as well as the connections between different pixels within it, thereby improving corn yield predictions.
@@
To simplify the learning process, we need to sample an appropriate number of pixels in each county. On the one hand, selecting fewer pixels might not capture the complete information in each county. On the other hand, over-sampling may result in an insufficient number of qualified counties, making it impossible to obtain sufficient data for training. Here, we randomly select 100 pixels from each county $i$, $i=1,...,n$, denoted as $\hat{P_{i}}=(P^{1}_{i},P^{2}_{i},...,P^{100}_{i})$. Our experiments in Section 4.4 have shown that these selected pixels will adequately capture the county’s information, as will be demonstrated in the experiment (Sec. 4.4). For simplicity of notation, we omit the subscript $i$ of $\hat{P_{i}}$ here, so it becomes $\hat{P}$. We aim to train a model $f$, $f(\hat{P})=y_{i}$, that can predict the yield accurately.
@@ 3.3 Attention mechanism implementation
The attention mechanism (Vaswani et al. 2017) in machine learning assigns varying importance to different input elements, enabling the model to focus on inputs that are more relevant to the output (Khan et al. 2022). This helps models perform better in complex tasks by learning context-specific representations. In our task, we not only use attention to implement MIL, but also to assign appropriate weights to the pixels in each county, hoping that attention can automatically determine the significance of each pixel. Ideally, it assigns higher weights for pure pixels and lower weights for mixed pixels. Suppose there are $100$ pixels in county $i$, $i=1,...,n$. For the feature $P_{i}^{j}$ of pixel $j$, $j=1,...,100$, our goal is to learn a weight $W_{i}^{j}$. We are aiming for such a model $f$, $f({W^{1}_{i}}^{\top}P^{1}_{i},{W^{2}_{i}}^{\top}P^{2}_{i},...,{W^{100}_{i}}^{\top}P^{100}_{i})=y_{i}$. In this way, we can reduce as much the noisy information in mixed pixels as possible.
@@
By combining the above equations, we denote the function of our attention module as $Att(\hat{P})$. Then our model $f$ becomes $f(Att(\hat{P}))=y_{i}$.
@@ 4.1.1 Performance Metrics
[…]
@@ 4.2 Model Performance Evaluation
$RMSE$ $R^{2}$ $RMSE$ $R^{2}$ $RMSE$ $R^{2}$ $RMSE$ $R^{2}$ $RMSE$ $R^{2}$
@@
$RMSE$ $R^{2}$ $RMSE$ $R^{2}$ $RMSE$ $R^{2}$ $RMSE$ $R^{2}$
@@
[…]
@@ 4.3 In-season yield prediction
Figure 7: In in-season county-level corn yield prediction in 2022: (a) $RMSE$ and (b) $R^{2}$.
@@ 4.4 The number of instances in a bag
Figure 8: The results of our model in 2022 under different numbers of selected pixels in one county: (a) $RMSE$ and (b) $R^{2}$.
@@ 5.1 Pixel mixing level calculation and preprocessing
To explore the role of attention, we first need to quantify the mixed pixels based on their definition, which is the proportion of the cornfield within a pixel of the feature dataset. Taking the MODIS dataset as an example, suppose one MODIS pixel contains $n$ CDL pixels. We represent each MODIS pixel as a binary list $S$ of length $n$, where each element $s_{i}$ represents a CDL pixel, $s_{i}\in\{0,1\}$ for $i=1,2,...,n$. Then, using GEE, we calculate the corn ratio, defined as the number of CDL pixels that represent corn within a MODIS pixel, to serve as the ground truth for our experiment:
@@
For the following analysis, the initial step of preprocessing is “bag normalization” by normalizing the corn ratio within each bag so that their sum equals one. Suppose we select $m$ pixels in each county $i$, $i=1,...,n$. In county $i$, we have a corn ratio $C_{i}^{j}$ for each pixel $j$, $j=1,...,m$. Then, the implementation of bag normalization is as follows:
@@ 5.3.1 Feature importance
where $value$ is the metric ($RMSE$ and $R^{2}$) of the feature being evaluated, $min$ is the metric of the least important feature (SOM), and $max$ is the metric of the most important feature (historical average yield). This formula scales the importance of each feature relative to the range between the minimum and maximum metric values.
@@ 5.5 Attention’s advantage over corn ratio
A simple solution to the mixed pixel problem is to directly calculate the corn ratio for each mixed pixel, and use it to weight each pixel before predicting corn yield. We conducted an experiment here to demonstrate that this method is not a preferable solution. We calculate the corn ratio using the CDL mask and the MODIS dataset, and then explicitly apply this corn ratio as a weight to the features of each pixel. Suppose we select $100$ pixels in each county $i$, $i=1,...,n$. We have the corn ratio $C_{i}^{j}$, feature vector $P_{i}^{j}$, and corn yield $y_{i}$ for each pixel $j$, $j=1,...,100$. Furthermore, the implementation of manual weighting is as follows:

### v4

@@ 2 Data acquisition
Category Variables Unit
Related
Properties
Spatial
Resolution
Source Latency
@@
Soil Organic Matter (SOM) $\mathrm{kg}$$\mathrm{m}^{2}$ Soil nutrient uptake
Cation Exchange Capacity (CEC) $\mathrm{cmol}$$\mathrm{kg}$
@@ 3.1 Problem definition
To formalize the county level crop yield prediction problem, let $n$ be the number of counties in the given study area. Let $I_{i}$ be the satellite imagery and $y_{i}$ be the yield for each county $i$, i=1,,ni=1,...,n. Each image is on the space $\mathbb{R}^{W\times H\times C}$, where $W$ and $H$ are the width and height of the image, and $C$ denotes the channels. We calculate VIs from the satellite imagery, along with other agricultural and environmental features (weather, soil properties, etc.), to form our county-level input feature $F_{i}$. From a machine learning perspective, our training objective is to develop a model $f$, $f(F_{i})=y_{i}$.
@@ 3.2 Multiple instance learning
Problem formulation for MIL in regression task For the county-level corn yield prediction task, we have the yield (label) of each county (bag), and each county is composed of a number of pixels (instances). By utilizing the detailed information of each county, MIL can improve prediction accuracy. Suppose there are $m_{i}$ pixels in each county $i$, i=1,,ni=1,...,n. To differentiate from the county-level feature $F$, we use $P$ to represent pixel-level features. For each pixel $j$, $j=1,...,m_{i}$, we extract its feature $P^{j}_{i}$. There exists a model $f$ with many-to-one mapping, namely $f(P^{1}_{i},P^{2}_{i},...\ldots,P^{m_{i}}_{i})=y_{i}$. We hope our model can automatically learn the structure of a county as well as the connections between different pixels within it, thereby improving corn yield predictions.
@@
To simplify the learning process, we need to sample an appropriate number of pixels in each county. On the one hand, selecting fewer pixels might not capture the complete information in each county. On the other hand, over-sampling may result in an insufficient number of qualified counties, making it impossible to obtain sufficient data for training. Here, we randomly select 100 pixels from each county $i$, i=1,,ni=1,...,n, denoted as $\hat{P_{i}}=(P^{1}_{i},P^{2}_{i},...,P^{100}_{i})$. Our experiments in Section 4.4 have shown that these selected pixels will adequately capture the county’s information, as will be demonstrated in the experiment (Sec. 4.4). For simplicity of notation, we omit the subscript $i$ of $\hat{P_{i}}$ here, so it becomes $\hat{P}$. We aim to train a model $f$, $f(\hat{P})=y_{i}$, that can predict the yield accurately.
@@ 3.3 Attention mechanism implementation
The attention mechanism (Vaswani et al. 2017) in machine learning assigns varying importance to different input elements, enabling the model to focus on inputs that are more relevant to the output (Khan et al. 2022). This helps models perform better in complex tasks by learning context-specific representations. In our task, we not only use attention to implement MIL, but also to assign appropriate weights to the pixels in each county, hoping that attention can automatically determine the significance of each pixel. Ideally, it assigns higher weights for pure pixels and lower weights for mixed pixels. Suppose there are $100$ pixels in county $i$, i=1,,ni=1,...,n. For the feature $P_{i}^{j}$ of pixel $j$, j=1,,100j=1,...,100, our goal is to learn a weight $W_{i}^{j}$. We are aiming for such a model $f$, $f({W^{1}_{i}}^{\top}P^{1}_{i},{W^{2}_{i}}^{\top}P^{2}_{i},...,{W^{100}_{i}}^{\top}P^{100}_{i})=y_{i}$. In this way, we can reduce as much the noisy information in mixed pixels as possible.
@@
By combining the above equations, we denote the function of our attention module as Att(P^)Att($\hat${P}). Then our model $f$ becomes $f(Att(\hat{P}))=y_{i}$.
@@ 4.1.1 Performance Metrics
where n is the total number of observations (counties or pixels). $y_{i}$ is the predicted value (predicted yield) for the $i\textsuperscript{th}$ observation (one pixel in RS imagery). $\hat{y}_{i}$ is the actual value (observed yield) for the $i\textsuperscript{th}$ observation. $\bar{y}$ is the mean of the observed data. The observed yield data are the official historical reports of the county level yields from USDA, NASS (USDA 2020b).
@@ 4.2 Model Performance Evaluation
RMSERMSE $R^{2}$ RMSERMSE $R^{2}$ RMSERMSE $R^{2}$ RMSERMSE $R^{2}$ RMSERMSE $R^{2}$
@@
RMSERMSE $R^{2}$ RMSERMSE $R^{2}$ RMSERMSE $R^{2}$ RMSERMSE $R^{2}$
@@
In comparison, the map corresponding to Ins-MIL exhibits slightly deeper colors, especially in Kansas in 2018 and in Iowa in 2020. This is because 2018 was particularly dry for many parts of Kansas, especially during key growth stages of corn (NASS 2018). The U.S. Drought Monitor observed parts of Kansas in severe to extreme drought conditions, particularly in the earlier part of the year (Drought.gov 2018). Furthermore, on August 10, 2020, a severe windstorm known as a “derecho” swept through Iowa (Service 2020). This storm brought extremely high winds, with some reports of wind speeds reaching over 100mphmph in certain areas. The derecho flattened cornfields, broke plants, and caused widespread destruction to crops across the central part of the state. This event had a significant negative impact on corn yields in Iowa for 2020. We use the error map to validate the connection between the storm and prediction error (Fig. 6). For cornfields affected by the storm, the satellite imagery does not change significantly in a short period of time. As a result, machine learning models tend to overestimate the yield of those affected areas. Indeed, in 2020, the results for this region show an overestimation, which is consistent with our hypothesis. The color intensities of these blocks in the maps of RF, LR, and RR are even deeper in Kansas and Iowa. Besides appearing darker in the two aforementioned regions, these maps also show extensive dark areas at the intersection of North Dakota, Minnesota, and South Dakota in 2021. This is because, in 2021, much of the upper Midwest, including North Dakota, South Dakota, and Minnesota, experienced severe drought conditions (Drought.gov 2021a, b, c; of Natural Resources 2021). In particular, North Dakota also experienced higher-than-average temperatures during the growing season (InForum 2021). Additionally, hot weather does not immediately affect satellite imagery, but it will impact the internal growth of plants and reduce yield. Therefore, high temperatures can also lead to corn yield overestimation, which is verified in our error map.
@@ 4.3 In-season yield prediction
Figure 7: In in-season county-level corn yield prediction in 2022: (a) RMSERMSE and (b) $R^{2}$.
@@ 4.4 The number of instances in a bag
Figure 8: The results of our model in 2022 under different numbers of selected pixels in one county: (a) RMSERMSE and (b) $R^{2}$.
@@ 5.1 Pixel mixing level calculation and preprocessing
To explore the role of attention, we first need to quantify the mixed pixels based on their definition, which is the proportion of the cornfield within a pixel of the feature dataset. Taking the MODIS dataset as an example, suppose one MODIS pixel contains $n$ CDL pixels. We represent each MODIS pixel as a binary list $S$ of length $n$, where each element $s_{i}$ represents a CDL pixel, $s_{i}\in\{0,1\}$ for i=1,2,,ni=1,2,...,n. Then, using GEE, we calculate the corn ratio, defined as the number of CDL pixels that represent corn within a MODIS pixel, to serve as the ground truth for our experiment:
@@
For the following analysis, the initial step of preprocessing is “bag normalization” by normalizing the corn ratio within each bag so that their sum equals one. Suppose we select $m$ pixels in each county $i$, i=1,,ni=1,...,n. In county $i$, we have a corn ratio $C_{i}^{j}$ for each pixel $j$, j=1,,mj=1,...,m. Then, the implementation of bag normalization is as follows:
@@ 5.3.1 Feature importance
where valuevalue is the metric ($RMSE$ and $R^{2}$) of the feature being evaluated, minmin is the metric of the least important feature (SOM), and maxmax is the metric of the most important feature (historical average yield). This formula scales the importance of each feature relative to the range between the minimum and maximum metric values.
@@ 5.5 Attention’s advantage over corn ratio
A simple solution to the mixed pixel problem is to directly calculate the corn ratio for each mixed pixel, and use it to weight each pixel before predicting corn yield. We conducted an experiment here to demonstrate that this method is not a preferable solution. We calculate the corn ratio using the CDL mask and the MODIS dataset, and then explicitly apply this corn ratio as a weight to the features of each pixel. Suppose we select $100$ pixels in each county $i$, i=1,,ni=1,...,n. We have the corn ratio $C_{i}^{j}$, feature vector $P_{i}^{j}$, and corn yield $y_{i}$ for each pixel $j$, j=1,,100j=1,...,100. Furthermore, the implementation of manual weighting is as follows: